### v1

# Expanding the Generative AI Design Space through Structured Prompting and Multimodal Interfaces

Abstract.
Text-based prompting remains the dominant interaction paradigm in generative AI, yet it often results in a high-friction experience for novice users, such as small business owners (SBOs), attempting to articulate creative or domain-specific goals for advertising. To investigate this challenge, we conducted a study with six SBOs in the United Kingdom, focusing on their advertising practices and perceptions and usage of AI tools in this context. Our findings surfaced two persistent breakdowns in current generative AI systems: first, the cognitive burden of prompt engineering, as users struggled to translate abstract creative goals into effective textual inputs; and second, the frequent generation of generic outputs that failed to align with users’ articulated brand vision. To address these issues, we developed ACAI (AI Co-Creation for Advertising and Inspiration), a multimodal, GenAI-powered advertisement creation tool designed to support novice designers by reimagining the prompt interface. ACAI features a structured, panel-based interface composed of three modules: the Branding Panel, the Audience & Goals Panel, and the Inspiration Board Panel to provide SBOs with outputs that align with their creative vision by reducing prompt ambiguity. This work contributes to HCI research on generative systems by showing how structured interfaces can foreground user-defined context to improve both alignment and promptability in novice workflows.
Structured Prompting, Interface Design, Novice Users, AI-Augmented Workflows, Multimodal Interfaces, GenAI
copyright: rightsretainedconference: ; April 27th, 2025; Yokohama, Japanjournalyear: 2025ccs: Human-centered computing User interface designccs: Human-centered computing Interaction designccs: Human-centered computing Empirical studies in HCI
1. Introduction
Small business owners (SBOs) represent the majority of business owners globally (Nations, 2024). Visual advertising and branding are critical to the success of small enterprises; yet, many SBOs lack formal training in design and often do not have access to professional creative services (Ballard, 2016; Bhalerao et al., 2022). The growing availability of generative AI systems presents new opportunities to support novice designers such as small business owners by reducing barriers to entry in advertising and branding and by broadening access to creative production tools.
To investigate how current generative systems support SBOs’ creative workflows, we conducted a formative study with six SBOs in Manchester, United Kingdom. Using semi-structured interviews, we examined participants’ advertising practices, experiences with generative AI, and expectations for AI-assisted design. To investigate how current generative systems support SBOs’ creative workflows, we conducted a formative study with six small business owners (SBOs) in Manchester, United Kingdom. Using semi-structured interviews, we examined participants’ advertising practices, experiences with generative AI, and expectations for AI-assisted design. Our findings surfaced three recurring challenges: (1) the cognitive effort required to translate abstract, domain-specific goals into usable prompts; (2) dissatisfaction with generic outputs that failed to reflect brand tone, values, or audience needs; and (3) concerns about losing creative control due to over-automation, signaling the need for systems that enable co-creation and preserve user agency. These themes align with ongoing HCI work on promptability (Morris, 2024) and the “gulf of envisioning” (Subramonyam et al., 2024), which highlight how non-expert users struggle to effectively guide generative systems.
To address this, we developed ACAI (AI Co-Creation for Advertising and Inspiration), a multimodal generative AI prototype that reimagines the prompt interface for novice designers. ACAI features a structured, panel-based user interface composed of three modules: the Branding & Asset Panel, the Audience & Goals Panel, and the Inspiration Board Panel. These modules enable users to specify brand assets, define campaign objectives, and upload visual references. A multimodal large language model (MLLM) processes these structured inputs into a unified “super prompt” that guides advertisement generation aligned with user-defined goals.
This work makes three primary contributions to HCI research on GenAI:
  • 1.
    A formative study surfacing promptability and brand-alignment challenges among SBOs using AI tools.
  • 2.
    Design requirements emphasizing multimodality and constraint-aware generation.
  • 3.
    The design of ACAI, a multimodal GenAI-powered prototype.
2. Related Work
2.1. Contextualization in Generative and Interactive Systems
Context-aware systems dynamically adapt to user needs by leveraging situational information captured through various modalities, such as GPS, sensors, and APIs (Dey, 2001; Chen, 2005; Aliannejadi et al., 2021; Clarizia et al., 2019). These systems have been widely adopted in domains such as mobile assistance and smart environments, where real-time responsiveness is essential. Building on this foundation, recent work has begun integrating context into generative workflows. For example, CARING-AI (Shi et al., 2025) generates spatially and temporally situated AR instructions using real-world environmental data, while ContextCam (Fan et al., 2024) incorporates user-captured context to guide image generation. These systems demonstrate how contextual signals can increase semantic alignment and enhance user engagement.
However, current text-to-image generation models primarily rely on text-based prompts and offer limited affordances for encoding structured user intent or domain-specific constraints (Lee et al., 2023). While these systems produce compelling outputs, they often fall short in supporting intention-aligned co-creation, particularly for novice users in task-specific domains. Motivated by these limitations, we position ACAI to bridge the gap between generic text-to-image systems and context-sensitive creative workflows by foregrounding explicit, user-defined contextualization.
2.2. Prompting Challenges and Interface Design for Novice Users
A growing body of HCI research has highlighted the barriers novice users face when interacting with prompt-based generative systems. Subramonyam et al.  (Subramonyam et al., 2024) identify this challenge as the gulf of envisioning, which refers to a cognitive disconnect between what users intend to create and how they are able to articulate those intentions in forms that a system can interpret. Similarly, Zamfirescu-Pereira et al. (Zamfirescu-Pereira et al., 2023) demonstrate that non-AI experts frequently default to ineffective prompting patterns, revealing a lack of support for articulating goals in ways that models can reliably interpret. To address these challenges, recent work has explored interaction designs that simplify how users articulate intent. PromptCrafter (Baek et al., 2023) introduces a dialogue-driven interface that helps users iteratively refine prompts with LLM support. These efforts reflect a broader shift in HCI toward designing promptable generative experiences, which are particularly effective in lowering entry barriers for novice and non-expert users. Building on this trajectory, our work foregrounds small business owners (SBOs) as a critical yet underrepresented user group in HCI. By designing for their situated creative workflows and contextual constraints, we contribute to ongoing efforts to develop task-aligned generative systems that broaden the inclusivity and usability of promptable interfaces for non-expert users.
Figure 1. Formative Study: 6 UK SBOs identify visual elements they would like to incorporate into their branding.
3. Formative Study
The last three authors conducted an in-person formative study with small business owners (SBOs) in Manchester, UK, to examine how SBOs perceive, engage with, and evaluate generative AI tools for advertising and brand development. The study aimed to understand existing branding practices, assess the alignment between generative AI value propositions and business needs, and identify adoption barriers. Participants were recruited through a professional research agency using an existing database of individuals open to participating in research. A screener questionnaire was distributed via email to identify eligible participants based on the following inclusion criteria: (1) over 18 years of age, (2) located in or near Manchester, and (3) a small business owner responsible for branding and the online presence of a business with a physical store or studio in the UK.
We conducted six in-person, semi-structured interviews, each lasting approximately 90 minutes. Sessions were audio recorded and transcribed for analysis. Participants received a £125 gift card as compensation. Businesses spanned diverse sectors, including food retail, entertainment, pet training, salon services, clothing retail, and personal services. Participants had varying levels of experience with generative AI tools, with most reporting some familiarity with platforms such as ChatGPT, Canva, or Gemini. Table 1 summarizes participant demographics and tool usage.
3.1. Key Findings from Formative Research
We analyzed the transcripts using reflexive thematic analysis. The following three findings emerged, informing the design of ACAI’s interface and interaction model:
  • 1.
    Challenges in Articulating Brand Intuition through Prompts: Participants struggled to articulate their brand vision using traditional text-based prompts. While they possessed a strong intuitive understanding of their brand identity, they found it difficult to externalize this understanding. One participant shared, “I instructed it to be funny, but that’s quite subjective, so I had to edit it quite heavily.” -P2. This finding underscores the tacit and iterative nature of branding, highlighting how limitations in promptability introduce friction into the creative workflows of novice users.
  • 2.
    Tensions Between Generativity and Control: While participants valued the efficiency offered by generative AI, they expressed concerns about losing creative control. Many voiced discomfort with over-automation and emphasized a preference for systems that facilitate co-creation, offering opportunities to adjust or override AI-generated outputs. One participant explained, “You couldn’t really change it, when AI produced that for you, there weren’t tweaks you could make to it… so that’s the limitation of it.” (P3). These tensions underscore the importance of interaction models that preserve user agency within AI-augmented workflows.
  • 3.
    Generic and Misaligned AI Outputs: Participants frequently described AI-generated advertisements as generic, repetitive, or misaligned with their brand’s tone and aesthetic. They noted that outputs often lacked the nuance necessary to reflect distinctive brand identities. This disconnect reinforces the need for generative systems that incorporate user-defined brand elements and contextual cues to produce brand-consistent outputs.
Gender Age AI Tools Business
Man 31–40 ChatGPT, Canva Food retail
Man 31–40 ChatGPT, Canva Entertainment
Man 18–40 ChatGPT, Gemini Pet training
Man 24–40 Gemini Salon
Woman 41–50 Canva Clothing retail
Man 41–50 ChatGPT, Canva Personal services
Table 1. Participant demographics and AI tool usage across business types
3.2. Design Requirements
Drawing on findings from our formative study, we outline design requirements to inform the development of generative AI systems that support small business owners in creating brand-aligned advertisements.
  • 1.
    DR1: Enable Multimodal Input and Interactive Style Selection for Creative Intent Expression
    Participants demonstrated an intuitive understanding of their brand identity and a sense of which visual elements aligned with their advertisements. However, they lacked the formal design vocabulary to articulate these elements through text prompts alone. Many participants found free-form prompting unintuitive, often requiring iterative refinements or abandonment due to inability to easily edit. This mismatch led to frustration when AI-generated content did not meet expectations. To mitigate this:
    • Multimodal Input Affordances: AI systems should allow users to specify creative intent through a combination of textual and visual modalities. Users should be able to upload or reference images alongside textual inputs, enabling a more comprehensive capture of their branding preferences.
    • Interactive Style Selection: Tools should support users in extracting, refining, and integrating design attributes, such as fonts, color schemes, and composition, from reference visuals into AI-generated outputs. This functionality should work in tandem with text prompts to enhance expressivity and alignment with users’ creative intent.
  • 2.
    DR2: Affordances and Constraints for Brand-Consistent AI Generation
    Participants frequently expressed concerns that AI-generated advertisements did not adequately reflect their brand’s established identity, often resulting in generic and misaligned outputs. To ensure brand fidelity, generative systems should incorporate both expressive affordances and protective constraints:
    • Brand Identity Encoding Affordances: AI tools should allow users to embed key elements of their brand identity, such as logos, specific fonts, color palettes, and tone of voice, into the generation pipeline. These affordances can mitigate users’ concerns of generic outputs.
    • Brand Consistency Constraints: The system should enforce constraints to preserve critical brand elements across outputs. These constraints act as fixed parameters (e.g., logo placement, business name inclusion, product-specific details) that ensure consistency while enabling flexibility across other design dimensions.
4. ACAI: AI Co-Creation for Advertising and Inspiration
Drawing on our formative study with small business owners (SBOs), we developed ACAI (AI Co-Creation for Advertising and Inspiration), a generative AI-powered prototype to support the creation of brand-aligned advertisements. ACAI is designed for users with limited design expertise but strong knowledge of their brand.
Figure 2. ACAI Architecture
4.1. System Overview
ACAI was developed as a standalone prototype to facilitate user testing sessions focused on co-creation practices among novice designers. The system leverages Gemini 1.5 Pro, a multimodal large language model (MLLM), to support its generative capabilities. ACAI’s architecture is organized into three core components: the User Input Layer, the Prompt Processing Layer, and the Output Generation Layer (see Figure 2).
  • 1.
    User Input Layer: The user input allows users to define brand attributes, advertising objectives, and stylistic preferences through three panels:
    • Branding & Asset Panel: Enables users to specify color palettes, typography, and brand values, as well as upload brand assets (logos, images, key visuals).
    • Audience & Goals Panel: Captures advertising objectives, target audience segmentation, and emotional tone preferences to ensure AI-generated ad brief aligns with marketing strategies.
    • Inspiration Board Panel: Allows users to upload reference images and interactively extract stylistic elements (e.g., fonts, elements, themes). A Multimodal Large Language Model (MLLM), such as Gemini, analyzes the images to extract regions of interest and generate structured textual descriptors for inclusion in the prompt.
    By organizing the prompt workflow into guided steps, ACAI supports novice users in articulating their intent more effectively, addressing common challenges in prompt formulation. (Subramonyam et al., 2024; Zamfirescu-Pereira et al., 2023).
  • 2.
    Processing Layer: In the processing layer, inputs from the three panels are synthesized into a “super prompt”, a structured, multimodal representation of the user’s design intent. This prompt includes both textual attributes (e.g., brand values, objectives) and image-derived descriptors (from the inspiration board). The resulting prompt is processed by the MLLM, which generates a textual Ad Brief designed to reflect the user’s brand identity and marketing goals.
  • 3.
    Output Generation Layer This layer incorporates the “Ad Brief Generator”, which converts AI-generated content into a structured brief. The generated Ad Brief contains the following three segments:
    • Summary: A high-level overview of the ad concept.
    • Background: Visual and contextual details.
    • Foreground: Key messaging and focal elements.
The Ad Brief functions as a guide for downstream visual production, whether by a designer or a future generative model. While ACAI does not currently produce visual assets directly, it facilitates human-AI co-creation by maintaining user control over content generation. Users may revise their input fields and regenerate the brief, supporting iterative cycles of ideation and refinement. Future iterations of the system will incorporate input weighting mechanisms to modulate the influence of specific parameters on the generated output.
Figure 3. ACAI Interface featuring the three main panels: Branding (A-E), Audience & Goals (F-H), and Inspiration Board (I-L) that structure user input for generating brand-aligned advertisement briefs.
4.2. Implementation walkthrough
To demonstrate how ACAI supports co-creation in real-world settings, we developed a demo brand named Arrivl, a travel agency managed by a hypothetical small business owner, Emily, based in London. Arrivl specializes in providing stress-free travel experiences, helping customers focus on their destination rather than the logistics of planning. To simulate authentic usage, we used a comprehensive set of branding assets for Arrivl, including a defined color palette, brand typography, brand values, logo, and three previously published advertisements. We also specified advertising objectives, target audience segmentation, and intended emotional tone to guide the AI-generated outputs. Additionally, we curated nine reference images for ACAI to analyze, enabling the system to extract relevant stylistic cues (Figure 3). The following section illustrates how ACAI assisted Emily in generating an AI-powered ad brief.
Emily begins by interacting with the Branding Panel, where she inputs key brand attributes that will shape the advertisement. She selects her brand’s colors in the color palette (Figure  3A), specifies Montserrat and Barlow as the preferred typefaces (Figure  3B), and uncertain about a preferred visual aesthetic (e.g., retro, natural, or minimalist), Emily chooses not to define a fixed Ad Style (Figure 3C). As a result, ACAI does not apply any stylistic constraints during initial Ad generation; however, the system remains flexible, allowing Emily to retroactively introduce stylistic preferences at any stage in the co-creation process. This design choice supports customization, even for users who may not have a fully articulated creative vision at the outset. Next, Emily proceeds to the Brand Values (Figure  3D). While she has the option to input a brand description manually, ACAI automatically generates an initial summary after she uploads brand assets (Figure  3E) and selects a target audience segment (Figure  3H). This autogenerated text synthesizes cues from her selected visuals and audience preferences, offering a system-generated interpretation of brand value. The text remains editable both before and after ad generation, enabling iterative refinement. This mechanism supports co-creation by lowering the initial barrier of articulating a brand narrative, particularly for novice users or those with limited experience expressing brand values in textual form. Emily then uploads and selects brand assets (Figure  3E), including Arrivl’s logo and three images from previous advertisements, although other types of visual material may also be used. While ACAI allows users to select between zero and four assets, Emily opts to include all four. These selected assets are treated as priority inputs and prominently featured in the generated ad brief, ensuring that user-provided content substantively informs the system’s output.
With Arrivl’s core branding traits set, Emily moves onto Audience & Goals Panel. Given Arrivl’s focus on travel, Emily enters ‘curious’ and ‘joyful’ as the expected user emotions (Figure 3F). Next, a range of advertisement goals (Figure 3G) are available for selection as well as a text field for a custom goal. Emily selects ‘brand awareness’ as the goal for both Arrivl’s primary and secondary audiences (Figure 3H). To explore potential market segments, Emily engages with the Target Audience (Figure 3H), inputting “Adventurer” and “The Mindful Traveler” as her primary and secondary segments. Based on her earlier inputs, including brand values and selected assets, ACAI generates associated demographic traits, interests, and motivations. This feature supports novice users in articulating audience strategy and allows for further refinement.
In the final stage, Emily interacts with the Inspiration Board (Figure 3I), where she engages with visual references to articulate her desired aesthetic. The interface displays a gallery of nine images (Figure 3I), which may include prior advertisements or aspirational references. Emily selects the top-left image from the gallery. Once selected, the image appears in a dedicated workspace beneath the gallery. ACAI’s underlying model, Gemini, analyzes the selected image and segments it into contextually salient visual components, such as the subject, background, and other prominent elements (in this case, a rock formation; (Figure 3J). Hovering over any segmented region reveals design-oriented annotations, which are natural language descriptions that articulate the stylistic or compositional qualities of the selected element. When Emily hovers over the rock formation, for example, the system surfaces the annotation: “weathered and covered in patches of moss, giving a sense of life and softness to its rough surface” (Figure 3K). These annotations assist novice designers like Emily in interpreting the aesthetic tone of the image and help inform decisions about which visual attributes to incorporate into the advertisement. To support finer-grained control, ACAI dynamically generates a set of adjustable sliders corresponding to image-specific features such as Depth of Field, Source of Light, Silhouette, Backdrop, and Adventure (Figure 3L).
5. Discussion
This paper contributes to ongoing HCI discourse on the promptability challenges faced by novice users in generative AI systems (Subramonyam et al., 2024; Zamfirescu-Pereira et al., 2023). Grounded in formative research with small business owners (SBOs) in the United Kingdom, we designed ACAI to support them in creating brand-aligned advertisements through structured prompting and multimodal interface interaction. Below, we reflect on ACAI’s broader implications for reimagining generative workflows and moving toward inclusive AI interface design.
5.1. Reimagining generative workflows
ACAI reimagines how novice users interact with generative AI systems by shifting from conventional text-based prompting to a structured, multimodal interaction paradigm. As a GenAI-powered advertisement creation tool, ACAI is designed to support small business owners in creating brand-aligned advertisements through an interface that reduces the cognitive burden of prompt engineering. By offering a more expressive and structured alternative to conventional text-based input paradigms, ACAI surfaces broader design considerations around ease of prompting and interaction flexibility that are central to evolving AI-augmented UI workflows.
While ACAI does not generate user interfaces per se, it contributes to the workshop’s central theme by demonstrating how AI functionality can inform interface architecture, input design, and creative tooling workflows. The system includes three structured input panels focused on branding, audience and goals, and inspiration, along with multimodal input mechanisms such as textual and image-based style selection. The Inspiration Board enables users to select reference images, which are analyzed by the system to extract design elements such as color, texture, and composition. When users hover over specific elements, the system generates descriptive design text, for example, ”rough textured, sun-drenched rock providing a natural platform and grounding element” and dynamically generates context-specific sliders to adjust stylistic and compositional attributes such as visual emphasis, mood, spatial depth, and aesthetic tone. These inputs guide content generation and demonstrate how UI workflows can be reimagined to better support AI-mediated creation.
5.2. Toward Inclusive AI Interface Design
In presenting this work, we aim to contribute a perspective on how designing AI-powered interfaces for novice creators can inform the development of more inclusive interfaces. We will highlight how ACAI’s structured and multimodal interface offers alternatives to conventional text-based prompting paradigms, and how such approaches can expand the design space for AI-enabled UI development. In doing so, we invite further reflection on how current generative AI design tooling might evolve beyond expert-oriented workflows toward more scaffolded and equitable interaction paradigms, which we believe are central to the themes this workshop aims to explore.
6. Conclusion and Future Work
In this paper, we presented ACAI (AI Co-Creation for Advertising and Inspiration), a multimodal generative AI prototype that reimagines the prompt interface for novice designers. Drawing on a formative study with small business owners, we identified key challenges in existing generative AI systems: the cognitive burden of prompt engineering and the generation of generic outputs that fail to align with users’ brand vision. ACAI addresses these challenges through a structured, panel-based interface that supports user input across brand identity, audience targeting, and visual style preferences.
Our approach demonstrates how structured interfaces can foreground user-defined context to improve both alignment and promptability in novice workflows. ACAI contributes to ongoing HCI research on co-creative tooling by offering an accessible, context-aware alternative to conventional prompt-based interaction  paradigms. We refer readers to (Karnatak et al., 2025) for a full discussion of the evaluation, which investigates how ACAI supports small business owners in creating brand-aligned advertisements within situated design workflows. Looking ahead, future research could explore how ACAI’s co-creation model might be adapted to other settings that require contextual grounding but involve users with limited design expertise.